The Southern Company (SO) has just paid an annual dividend of $3 per share. If the expected growth rate for Southern Co. is 10% and your required rated of return is 16%, how much are you willing to pay for this stock?

1)$55

2)$50

3)$46.50

4)$52.50
Answer #1

Constant growth dividend discount model

Price = \frac{Div_{1}}{r - g}

Div1 = 3 * (1 + 0.1) = $3.3

Price = \frac{3.3}{0.16 - 0.1}

Price = \$55

We are willing to pay $55 for this stock.

Can you please upvote? Thank You :-)